# Verti-Bench: A General and Scalable Off-Road Mobility Benchmark for Vertically Challenging Terrain

### v1

@@ Abstract
[…]
@@ I Introduction
In this work, we present Verti-Bench (Fig. LABEL:fig:cover), a general and scalable off-road mobility benchmark that focuses on extremely rugged, vertically challenging terrain with a variety of unstructured off-road features. Based on a high-fidelity multi-physics dynamics simulator, Chrono [28], Verti-Bench encapsulates variations in four orthogonal dimensions: Using the Sliced Wasserstein Autoencoder (SWAE) [29] and real-world off-road terrain data, off-road geometry is represented as a diverse set of 2.5D elevation maps; Ten terrain semantics classes, including seven rigid and three deformable, are designed with different distributions of physics parameters, e.g., friction coefficient, cohesive effect, and soil stiffness; Different types of natural obstacles, e.g., boulders and vegetations, are randomly distributed based on different densities; A set of off-road vehicles with different scales (from 1/10th to full scale) and actuation mechanisms (4-, 6-, and 8-wheeled and 2-tracked chassis, single- and double-wishbone, multilink, toebar leaf-spring, and special tensioning suspensions, as well as pitman-arm, rack-and-pinion, toebar, bellcrank/rotary arm, and differential steering) are provided, with the possibility of adding customized vehicles. Using Verti-Bench, we also provide datasets from expert demonstration, random exploration, failure cases (rolling over and getting stuck), as well as a gym-like interface for Reinforcement Learning (RL). We use Verti-Bench to benchmark ten off-road mobility systems, present our findings, and identify future off-road mobility research directions. To summarize, our contributions are:
@@ III-F Metrics
Verti-Bench automatically computes a set of standard metrics to quantify off-road mobility performance, while additional metrics can be customized as needed: Success Rate captures the percentage of successful trials over total number of attempts; Traversal Time indicates how long it takes to finish a successful traversal; Roll and Pitch describes the stability of the vehicle during traversal, whose raw and absolute values can be used to derive mean, variance, and maximum; Vehicle actions, such as throttle and steering, can be used as metrics to quantify energy consumption, planning confidence, and path smoothness, along with other metrics. Verti-Bench provides infrastructure to save raw vehicle-terrain interaction data as well as to compute performance metrics.
@@ III-G3 Failure Cases
To enable future data-driven off-road mobility by preventing vehicle failures, we also curate a dataset of failure cases by providing the last ten seconds of trajectory before the vehicle rolls over or gets stuck. The failure cases can be used to learn high-cost regions to be avoided in data-driven mobility systems.
@@ III-G4 RL Interface
Although not being a main purpose of Verti-Bench, we also provide a gym-like RL interface so that vehicles can learn off-road mobility through trial-and-error experiences in Verti-Bench. Vehicle states and actions, as well as reward functions, can be customized by our interface, which communicates with existing RL algorithm implementations.
@@ IV Evaluation and Discussions
Figure 4: Success Rate, Traversal Time, Roll, and Pitch of Ten Off-Road Mobility Systems on 1000 Navigation Tasks.
@@ IV-A Off-Road Mobility Systems for Evaluation
We reach out to the authors of the original papers for their implementations of their mobility systems. Considering Verti-Bench’s focus on off-road mobility evaluation, we make minimal modifications to their implementations to interface with Verti-Bench so that their mobility systems are no longer dependent on any perception system. For example, visual odometry inputs are replaced with ground truth vehicle states from Verti-Bench; Real-world elevation mapping systems are skipped by directly providing their systems with ground truth Verti-Bench elevation maps.
@@ IV-B Evaluation Results and Discussions
The evaluation results are shown in Fig. 4, including percentage of Succuss Rate and mean and variance of Traversal Time, Roll, and Pitch. We also divide the evaluation results with respect to the three elevation levels.
@@
In general, navigation performance significantly declines with increasing elevation levels, including reduced Success Rate and increased Traversal Time, Roll, and Pitch. In addition to mean, the variance of Roll and Pitch also drastically increases, indicating much less stable vehicle chassis when traversing high elevation environments. Among the three categories, end-to-end learned mobility systems achieve the worst performance, while hybrid systems outperform the other two in general. While ACL is expected to outperform MCL, the results suggest otherwise. Such results indicate that end-to-end learning methods, trained from other sources, still have much room for improvement in terms of generalization in Verti-Bench. For hybrid systems, TAL and TNT are the two top performing planners among all systems overall, achieving the highest Success Rate in all cases and lowest Roll and Pitch angle in most cases. WMVCT and MPPI-6 achieve good Success Rate in high elevation environments, but with large Roll and Pitch. Classical planners perform in between their end-to-end and hybrid counterparts. PID, due to its simplicity and robustness, performs very well in low elevation environments, with EH catching up on Success Rate when facing higher elevation. MPPI does not perform well in most cases and only outperforms PID in terms of Success Rate in high elevation environments. Notice that Traversal Time is only averaged over successful trials and thus only indicates how fast a mobility system is given navigation success.
@@
Our evaluation results indicate the potential of hybrid mobility systems to tackle vertically challenging terrain by combining the best of both worlds of classical and learning approaches. The overall success of TAL and TNT indicates the importance of an accurate 6-DoF kinodynamic model enabled by sophisticated learning techniques in conjunction with a sampling-based motion planner. MPPI-6, with a 6-DoF kinodynamic model based on a simplistic neural network, underperforms TAL and TNT, while the inaccuracies introduced by WMVCT’s efficient 6-DoF decomposition lead to the worst mobility performance among hybrid systems. The degraded performance of all hybrid systems in high elevation environments motivates further research, potentially to both increase the kinodynamic modeling accuracy and improve the sampling-based motion planner. On the other hand, it is surprising to see the superior performance of the simple PID planner in low elevation environments compared to the more sophisticated EH, whose advantage only starts to slightly emerge in high elevation environments. This observation reveals a tradeoff between system complexity and performance when facing simple environments. One potential future research direction is to develop off-road mobility systems composed of multiple planners with different complexities and specialties to fit different environments [56]. Lastly, research of end-to-end learning approaches, despite their recent success in relatively benign indoor or on-road enviornments, still needs to focus on robustly generalizing to out-of-distribution scenarios, which are very common to encounter in off-road enviornments.
@@ V Real-World Validation
Figure 5: Physical Off-Road Testbed Similar to Verti-Bench.
Table I shows the physical validation results of the three mobility systems on both low and high elevation testbeds. In general, the trend of the physical experiment results matches with that of Verti-Bench evaluation results: On the low elevation testbed, both PID and TNT are able to finish all five trials, while ACL still suffers from poor generalization. PID is still the fastest due to a lack of consideration of elevation. The roll and pitch angles are small and stable for all systems due to the lower and smoother terrain elevation; On the high elevation testbed, the difference between TNT and PID starts to emerge. TNT succeeds all five trials, while PID fails two. TNT and PID also exhibit smaller roll and pitch angle respectively. PID is still the fastest. ACL, similar to all previous cases, fails three trials and experiences the largest roll and pitch angles. Notice that due to the difficulty in conducting physical experiments, we only limit to physically evaluating three systems and five trials each, totaling 30 trials. This contrast against our 10000 trials (ten systems, 1000 trials each) in Verti-Bench, which can achieve much better statistical significance, further suggests the utility of Verti-Bench to evaluate off-road mobility.
TABLE I: Physical Validation of PID, ACL, and TNT: Success Rate, Traversal Time, Roll, and Pitch.

### v2

@@ Abstract
Recent advancement in off-road autonomy has shown promises in deploying autonomous mobile robots in outdoor off-road environments. Encouraging results have been reported from both simulated and real-world experiments. However, unlike evaluating off-road perception tasks on static datasets, benchmarking off-road mobility still faces significant challenges due to a variety of factors, including variations in vehicle platforms and terrain properties. Furthermore, different vehicle-terrain interactions need to be unfolded during mobility evaluation, which requires the mobility systems to interact with the environments instead of comparing against a pre-collected dataset. In this paper, we present Verti-Bench, a mobility benchmark that focuses on extremely rugged, vertically challenging off-road environments. 100 unique off-road environments and 1000 distinct navigation tasks with millions of off-road terrain properties, including a variety of geometry and semantics, rigid and deformable surfaces, and large natural obstacles, provide standardized and objective evaluation in high-fidelity multi-physics simulation. Verti-Bench is also scalable to various vehicle platforms with different scales and actuation mechanisms. We also provide datasets from expert demonstration, random exploration, failure cases (rolling over and getting stuck), as well as a gym-like interface for reinforcement learning. We use Verti-Bench to benchmark ten off-road mobility systems, present our findings, and identify future off-road mobility research directions. Verti-Bench project website can be found at https://cs.gmu.edu/~xiao/Research/Verti-Bench.
@@ I Introduction
In this work, we present Verti-Bench (Fig. LABEL:fig:cover), a general and scalable off-road mobility benchmark that focuses on extremely rugged, vertically challenging terrain with a variety of unstructured off-road features. The main goal of Verti-Bench is to directly compare the performances of different off-road mobility systems. Based on a high-fidelity multi-physics dynamics simulator, Chrono [28], Verti-Bench encapsulates variations in four orthogonal dimensions: Using the Sliced Wasserstein Autoencoder (SWAE) [29] and real-world off-road terrain data, off-road geometry is represented as a diverse set of 2.5D elevation maps; Ten terrain semantics classes, including seven rigid and three deformable, are designed with different distributions of physics parameters, e.g., friction coefficient, cohesive effect, and soil stiffness; Different types of natural obstacles, e.g., boulders and vegetations, are randomly distributed based on different densities; A set of off-road vehicles with different scales (from 1/10th to full scale) and actuation mechanisms (4-, 6-, and 8-wheeled and 2-tracked chassis, single- and double-wishbone, multilink, toebar leaf-spring, and special tensioning suspensions, as well as pitman-arm, rack-and-pinion, toebar, bellcrank/rotary arm, and differential steering) are provided, with the possibility of adding customized vehicles. Using Verti-Bench, we also provide datasets from expert demonstration, random exploration, failure cases (rolling over and getting stuck), as well as a gym-like interface for Reinforcement Learning (RL). We use Verti-Bench to benchmark ten off-road mobility systems, present our findings, and identify future off-road mobility research directions. To summarize, our contributions are:
@@ III-F Metrics
Verti-Bench automatically computes a set of standard metrics to quantify off-road mobility performance, while additional metrics can be customized as needed: A successful completion is defined as the robot reaching within 10 meters of the goal in less than 60 seconds. A successful trial may include contacts with obstacles. Success Rate captures the percentage of successful trials over total number of attempts; Traversal Time indicates how long it takes to finish a successful traversal; Roll and Pitch describes the stability of the vehicle during traversal, whose raw and absolute values can be used to derive mean, variance, and maximum; Vehicle actions, such as throttle and steering, can be used as metrics to quantify energy consumption, planning confidence, and path smoothness, along with other metrics. Verti-Bench provides infrastructure to save raw vehicle-terrain interaction data as well as to compute performance metrics.
@@ III-G3 Failure Cases
To enable future data-driven off-road mobility by preventing vehicle failures, we also curate a dataset of failure cases by providing the last ten seconds of trajectory before the vehicle rolls over or gets stuck. We divide all failure cases into two categories: (1) Rollover: the robot has a $>30$° roll angle at the end of a failed trial; (2) Stuck: the robot does not move more than 1 m during the last 10 seconds of a failed trial (with $\leq 30$° roll). The failure cases can be used to learn high-cost regions to be avoided in data-driven mobility systems.
@@ III-G4 RL Interface
Although not being a main purpose of Verti-Bench, we also provide a gym-like RL interface so that vehicles can learn off-road mobility through trial-and-error experiences in Verti-Bench. Vehicle states and actions, as well as reward functions, can be customized by our interface, which communicates with existing RL algorithm implementations. Verti-Bench provides interfaces of different inputs to the mobility systems, such as elevation and semantic maps, robot states, and raw sensor data. Users can choose appropriate inputs based on their evaluation needs.
@@ IV Evaluation and Discussions
Figure 4: Success Rate, Traversal Time, Roll, and Pitch with respect to Elevation Level (top), Terrain Semantics (middle), and Obstacle Density (bottom) of Ten Off-Road Mobility Systems on 1000 Navigation Tasks.
@@ IV-A Off-Road Mobility Systems for Evaluation
We reach out to the authors of the original papers for their implementations of their mobility systems. Considering Verti-Bench’s focus on off-road mobility evaluation, we make minimal modifications to their implementations to interface with Verti-Bench so that their mobility systems are no longer dependent on any perception system. For example, visual odometry inputs are replaced with ground truth vehicle states from Verti-Bench; Real-world elevation mapping systems are skipped by directly providing their systems with ground truth Verti-Bench elevation maps; All learning systems and components in the evaluation are not trained in Verti-Bench.
@@ IV-B Evaluation Results and Discussions
The evaluation results are shown in Fig. 4, including percentage of Succuss Rate and mean and variance of Traversal Time, Roll, and Pitch with respect to three types of elevation level, terrain semantics, and obstacle density for all systems and tasks. We also present an analysis of failure cases shown in Table I. All 1000 Verti-Bench navigation tasks have been used to evaluate each mobility system (no task has been used for training). A complete evaluation of each system requires approximately 10 hours. All task configurations have been documented in YAML files, which can be used by external research teams to replicate and expand these evaluations. Additionally, we provide our terrain generation pipeline, allowing researchers to extend or customize environmental parameters according to their specific research objectives and experiment requirements.
TABLE I: Failure Case Analysis of 1000 tasks with Ten Mobility Systems
PID EH MPPI RL MCL ACL WMVCT MPPI-6 TAL TNT
Rollover (%) 14.7 16.3 17.9 14.3 16.9 15.4 18.9 18.0 17.0 17.5
Stuck (%) 36.5 39.4 40.9 75.7 41.3 48.0 39.6 35.0 24.6 25.1
Success (%) 48.8 44.3 41.2 10.0 41.8 36.6 41.5 47.0 58.4 57.4
@@
In general, navigation performance significantly declines with increasing elevation levels, deformable surfaces, and obstacle densities, including reduced Success Rate and increased Traversal Time, Roll, and Pitch. In addition to mean, the variance of Roll and Pitch also drastically increases, indicating much less stable vehicle chassis when traversing high elevation, deformable, and obstacle-dense environments. Among the three categories, end-to-end learned mobility systems achieve the worst performance, while hybrid systems outperform the other two in general. While ACL is expected to outperform MCL, the results suggest otherwise. Such results indicate that end-to-end learning methods, trained from other sources, still have much room for improvement in terms of generalization in Verti-Bench. Considering failure cases, the systems fail more frequently due to getting stuck than rolling over. While the failure rates due to rollover are relatively consistent across all systems, failure rates due to getting stuck differ significantly, with RL getting stuck 75.7% of time, compared to TAL and TNT with the lowest getting-stuck rates (24.6% and 25.1% respectively).
In terms of elevation, for hybrid systems, TAL and TNT are the two top performing planners among all systems overall, achieving the highest Success Rate in all cases and lowest Roll and Pitch angle in most cases. WMVCT and MPPI-6 achieve good Success Rate in high elevation environments, but with large Roll and Pitch. Classical planners perform in between their end-to-end and hybrid counterparts. PID, due to its simplicity and robustness, performs very well in low elevation environments, with EH catching up on Success Rate when facing higher elevation. MPPI does not perform well in most cases and only outperforms PID in terms of Success Rate in high elevation environments. Notice that Traversal Time is only averaged over successful trials and thus only indicates how fast a mobility system is given navigation success.
For terrain semantics, all systems perform best on rigid terrain, with TAL and TNT achieving highest Success Rates above 60%. Performance drops significantly on deformable surfaces, with even the highest only achieving around 40% Success Rates while end-to-end systems struggle below 10%. Mixed terrain results fall between rigid and deformable terrain. Deformable surfaces also show increased roll and pitch variance across all systems, indicating less stable navigation and highlighting the challenge of modeling vehicle dynamics on unpredictable surfaces.
As obstacle density increases, Success Rate declines and Traversal Time increases across all systems, with the performance gap between hybrid and other systems widening in dense environments. Obstacle density does not directly affect vehicle stability, showing similar Roll and Pitch mean and variance.
@@
Our evaluation results indicate the potential of hybrid mobility systems to tackle vertically challenging terrain by combining the best of both worlds of classical and learning approaches. The overall success of TAL and TNT indicates the importance of an accurate 6-DoF kinodynamic model enabled by sophisticated learning techniques in conjunction with a sampling-based motion planner. MPPI-6, with a 6-DoF kinodynamic model based on a simplistic neural network, underperforms TAL and TNT, while the inaccuracies introduced by WMVCT’s efficient 6-DoF decomposition lead to the worst mobility performance among hybrid systems. The degraded performance of all hybrid systems when facing high elevation, deformable surfaces, and dense obstacles motivates further research, potentially to both increase the kinodynamic modeling accuracy and improve the sampling-based motion planner. On the other hand, it is surprising to see the superior performance of the simple PID planner in low elevation environments compared to the more sophisticated EH, whose advantage only starts to slightly emerge in high elevation environments. This observation reveals a tradeoff between system complexity and performance when facing simple environments. One potential future research direction is to develop off-road mobility systems composed of multiple planners with different complexities and specialties to fit different environments [56]. Lastly, research of end-to-end learning approaches, despite their recent success in relatively benign indoor or on-road enviornments, still needs to focus on robustly generalizing to out-of-distribution scenarios, which are very common to encounter in off-road environments.
@@ V Real-World Validation
Figure 5: Physical Off-Road Testbed Similar to Verti-Bench.
@@
Table II shows the physical validation results of the three mobility systems on both low and high elevation testbeds. In general, the trend of the physical experiment results matches with that of Verti-Bench evaluation results: On the low elevation testbed, both PID and TNT are able to finish all five trials, while ACL still suffers from poor generalization. PID is still the fastest due to a lack of consideration of elevation. The roll and pitch angles are small and stable for all systems due to the lower and smoother terrain elevation; On the high elevation testbed, the difference between TNT and PID starts to emerge. TNT succeeds all five trials, while PID fails two. TNT and PID also exhibit smaller roll and pitch angle respectively. PID is still the fastest. ACL, similar to all previous cases, fails three trials and experiences the largest roll and pitch angles. Notice that due to the difficulty in conducting physical experiments, we only limit to physically evaluating three systems and five trials each, totaling 30 trials. This contrast against our 10000 trials (ten systems, 1000 trials each) in Verti-Bench, which can achieve much better statistical significance, further suggests the utility of Verti-Bench to evaluate off-road mobility.
TABLE II: Physical Validation of PID, ACL, and TNT: Success Rate, Traversal Time, Roll, and Pitch.
@@ VII Conclusions
Acknowledgments
This work has taken place in the RobotiXX Laboratory at George Mason University. RobotiXX research is supported by National Science Foundation (NSF, 2350352), Army Research Office (ARO, W911NF2320004, W911NF2420027, W911NF2520011), Air Force Research Laboratory (AFRL), US Air Forces Central (AFCENT), Google DeepMind (GDM), Clearpath Robotics, Raytheon Technologies (RTX), Tangenta, Mason Innovation Exchange (MIX), and Walmart.